1. A centrifuge starts from rest and reaches an angular velocity of 30000 rpm in 5s. What is the angular acceleration(in rad/s2) of this device?

2. A flywheel rotating at 20 rpm undergoes an angular acceleration of 2.0 rad/s2 for precisely 5 revolutions. What is the angular velocity of the flywheel at the end of this acceleration?
Answer #1

Angular acceleration

= final angular velocity- initial angular velocity / time

= 18.7 - 75.4/ 10= - 5.67 rad/ s^2

Angular displacement = wt + 0.5 x angular acceleration × t^2

= 75.4 x 10 - 0.5 x 5.67 x 10 x 10 = 470.5 rad

Linear displacement = angular displacement x radius

= 470.5 x 0.356/2= 83.749 m